ashley buys 3 items that cost d dollars each when she pays the cashier for the items she receives $8 in change write an expression that represents the amount of money ashley gave the cashier
Answer by jim_thompson5910(35256)

3 items cost d dollars each. In total, they cost 3d dollars

Let's cay she pays c dollars in cash.

The amount of change is the difference in the two: c - 3d

we know this change is $8, therefore, c-3d = 8

Isolate c:

c-3d = 8

c-3d+3d = 8+3d

c = 3d + 8

The final answer is 3d + 8

Side note: we have NOT factored in taxes of any kind